OpenStudy (anonymous):

Integral of (5x^2)/((x^2-4)(3x^2+8)). How do you work it with the difference of squares in there?

OpenStudy (anonymous):

use equation editor... \[\int\limits \frac{ 5x^2 }{ \left( x^2-4 \right) \left( 3x^2+8 \right) }\,dx\]like this?

OpenStudy (anonymous):

you have to use partial fraction decomposition...

myininaya (myininaya):

you can factor a difference of squares a^2-b^2=(a-b)(a+b)

OpenStudy (anonymous):

i would hope that @kreimer20 knows that... this is calculus after all!

myininaya (myininaya):

i know its calculus

OpenStudy (anonymous):

i know

myininaya (myininaya):

he just asked how you do it with a difference of squares that particular difference of squares can be factored as a linear * linear

myininaya (myininaya):

when writing as partial fractions you would do A/linear +B/the other linear +(Cx+D)/the quadratic you have there

OpenStudy (anonymous):

yes, and as student studying calculus, she or he should know how to factor a difference of squares

myininaya (myininaya):

some people forget algebra just reminding him just in case

myininaya (myininaya):

i have seen students do some crazy algebra things in calculus

myininaya (myininaya):

like as in crazy wrong

OpenStudy (anonymous):

and unfortunately, some never learned. me, too! it's so funny (sic) that people struggle with calculus because they don't know their algebra or trig!

myininaya (myininaya):

i'm sure there was a couple of things i forgot when i was taking calculus (i think :p)

zepdrix (zepdrix):

I always found Difference of Cubes formula hard to remember :) lol

OpenStudy (anonymous):

also, see Paul's online math notes (here's a link). The site does a great job of working through the partial fraction decomposition. The attachments I sent previously are also a great resource! http://tutorial.math.lamar.edu/Classes/CalcII/PartialFractions.aspx

myininaya (myininaya):

at one point i was like i'm don't going to learn that formula i can find one zero for a difference or sum of cubes then use that to find the other factor by using division

OpenStudy (anonymous):

I know how to factor the difference of squares... Just forgot that I could do that to this problem. My professor has made it a little confusing for me...

OpenStudy (anonymous):

sum and difference of cubes is easy if you remember to repeat with out cubes, and then perfect square trinomial with different sign and no doubling of middle term... \[\left( a^3+b^3 \right)=\left( a+b \right)\left( a^2-ab+b^2 \right)\] \[\left( a^3-b^3 \right)=\left( a-b \right)\left( a^2+ab+b^2 \right)\]

myininaya (myininaya):

yeah at one point i know those by heart now seen them for like 8000 years

myininaya (myininaya):

so @kreimer20 i guess you can go further on this problem now? let me or us know if you need further help.

OpenStudy (anonymous):

so, @kreimer20 are you good? hold on to those pdfs and the link to Paul's online math notes.